if the sum of the measures of the interior angles of a polygon equals the sum of the measures of its exterior angles, how many sides does it have?
Answer by Gogonati(855)

In each polygon the sum of exterior angels is 360 degree.
Let n the number of interior angels of the polygon. The sum of interior angels is:
(n-2)180. Since the measure of interior angels is equal to the measure of
exterior angels we write:(n-2)180=360 => n-2=360/180 => n-2=2 => n=4.
Answer:The requested polygon is the square.